### v2

# RDF Stream Taxonomy: Systematizing RDF Stream Types in Research and Practice

Abstract
Over the years, RDF streaming was explored in research and practice from many angles, resulting in a wide range of RDF stream definitions. This variety presents a major challenge in discussing and integrating streaming solutions, due to the lack of a common language. This work attempts to address this critical research gap, by systematizing RDF stream types present in the literature in a novel taxonomy. The proposed RDF Stream Taxonomy (RDF-STaX) is embodied in an OWL 2 DL ontology that follows the FAIR principles, making it readily applicable in practice. Extensive documentation and additional resources are provided, to foster the adoption of the ontology. Two realized use cases are presented, demonstrating the usefulness of the resource in discussing research works and annotating streaming datasets. Another result of this contribution is the novel nanopublications dataset, which serves as a collaborative, living state-of-the-art review of RDF streaming. The aim of RDF-STaX is to address a real need of the community for a better way to systematize and describe RDF streams. The resource is designed to help drive innovation in RDF streaming, by fostering scientific discussion, cooperation, and tool interoperability.
Keywords:
RDF stream Taxonomy Ontology Nanopublications.
Resource type: Ontology, documentation, nanopublications dataset
DOI: 10.5281/zenodo.10072907
URL: https://w3id.org/stax/
1 Introduction
Over the years, the concept of RDF streams was investigated from many angles, resulting in streaming serializations, querying and reasoning engines, programming frameworks, datasets, and other resources [5, 50, 58, 60]. Without a doubt, streaming RDF solutions are in demand, and more innovation is still needed. However, the landscape of RDF streaming terminology is fragmented, with many dissimilar streaming task formulations and incompatible solutions.
The term “RDF stream” is used in the literature to refer to many different concepts, ranging from theoretical models [11], to very technical ones, focusing on the representation and practical usage [17, 45, 50]. As noted in past research [14, 35, 49], RDF streaming systems need more standardization in stream exchange protocols and stream task definitions. At the same time, if the research is to solve real-world problems, the variety of the existing RDF streaming task formulations should be embraced, not limiting oneself to one specific type of RDF streams. However, up until now, there was very little academic discussion about the existing RDF stream types – what stream types are used, how they relate to each other, and which use cases are they applicable to. This paper aims to address this critical research gap.
The main contribution of this work is the RDF Stream Taxonomy (RDF-STaX), embodied in an OWL 2 DL ontology. This novel taxonomy systematizes the existing RDF stream types on the basis of a broad state of the art review. RDF-STaX aims to be a practical and useful resource for researchers and practitioners, enabling them to semantically annotate RDF stream types in publications, software, and streams published on the Web. To achieve this, RDF-STaX is accompanied with extensive documentation, including usage examples. Finally, two implemented applications of RDF-STaX are presented – a living state-of-the-art review of RDF streaming using nanopublications, and annotating stream types in RiverBench datasets.
2 Literature Review
The primary goal of this paper is to introduce a taxonomy of RDF stream types, which naturally must be grounded in the state of the art to be relevant. As no comprehensive review of this topic is available, here we attempt to summarize the most significant works that use the concept of RDF streams.
The scope of this study only includes RDF streams that can be fully represented using the RDF 1.1 W3C Recommendation [10]. This notably excludes stream models that do not have a direct translation to RDF, such as RSP-QL [11]. However, due to the importance of such models, they are briefly discussed below. The reason behind this choice is that only streams fully representable in RDF can be easily exchanged and published on the Web, while the theoretical models have more limited applicability, e.g., only within one specific solution (e.g., a stream reasoning engine). As shown below, this scope choice still leaves a lot of variety to be explored.
The review was conducted by starting from several survey papers and books that focus on areas of research related to RDF streaming [5, 29, 39, 40, 43, 46, 53, 58, 60, 64], such as RDF data management, linked streaming data, and stream reasoning. Besides the works mentioned in these reviews, additional papers and software were added based on the authors’ expertise and by searching for keywords “RDF stream” and “RDF streaming” in Google Scholar. The review focused on works which use streams of RDF-based elements (triples, quads, graphs, datasets). For many of the examined works, determining the exact nature of the stream was not trivial, due to unclear descriptions and a lack of standardized terminology. Therefore, not only the papers were examined, but also the code and documentation, where available.
2.1 RDF Streaming Protocols
One area of research that mentions RDF streams are works focusing on streaming protocols for transmitting RDF data. One of the earlier examples is Streaming HDT (S-HDT) [26], which facilitates streaming RDF data in IoT devices. More specifically, producers (IoT devices) emit a stream of compressed, unnamed RDF graphs. RDSZ [18] is a later method that also attempts streaming compressed RDF graphs over the network. ERI [17] is a similar approach, however, it has several views of the RDF stream. At the highest level, ERI streams a continuous sequence of triples that is then divided into discrete blocks (RDF graphs). Then, on the lowest level, these graphs are decomposed into subject-molecules (graphs with a single RDF subject) that are then streamed over the network. Therefore, ERI has three different views of the streaming problem, with each corresponding to a different level in the protocol.
The RDF EXI communication protocol for constrained devices [28] focuses on streaming RDF graphs. Finally, the recent Jelly protocol [50] treats RDF streams similarly to ERI, with it streaming a sequence of RDF statements divided into discrete elements (stream frames). The main difference is that Jelly can also stream quad statements, making it generalizable to RDF datasets.
2.2 Stream Processing and Reasoning
There is a large body of research on the topics of RDF stream reasoning and stream processing, where a special focus is often placed on the time annotations of stream elements. We could find one work in this category that did not consider the temporal aspect, an early proposal for a streaming SPARQL engine from 2007 by Groppe et al. [23]. The proposed engine executes queries over streams of RDF triples.
Most publications in this category focus on the inner workings of querying/reasoning engines, and adopt task formulations that are not fully grounded in the RDF specification. The most popular such formulation is the stream of timestamped triples [1, 3, 4, 8, 11, 13, 30, 34, 57, 64], where a stream is a sequence of $\langle triple,timestamp\rangle$ tuples. There are many variations of this task, with time intervals instead of timestamps [1], multiple timestamps per triple [12, 57], and different time domains. However, this model (and other equivalent or similar models, e.g., time-varying graphs [14, 33, 57]) by itself is not enough to fully realize an RDF stream – this would require encoding the time information in RDF, which is often omitted in these works.
The lack of grounding in the RDF specification of the timestamped triple model is a known problem [8], and solutions to it were proposed. The most prominent formulation is the stream of timestamped named graphs, where each stream element is a named RDF graph. The timestamp (or other metadata) of this graph is added by making statements in the default graph about the named graph’s name node [7, 41, 63]. This representation was possibly first proposed by Tappolet & Bernstein in 2009 [56], and was then expanded by the W3C RSP Community Group, which created a draft specification for the RSP Data Model [47]. The RSP Data Model uses the aforementioned formulation to define RDF streams, but also mentions that RSP engines may consume streams of graphs, datasets, or named graphs and then turn them into RDF streams proper by associating each element with a timestamp. The RSP Data Model is the most mature and best-grounded model for representing timestamped streams, having been drafted by the community, and even proposing the formal semantics of RDF streams. However, it was never made into a W3C Recommendation, and the RSP Community Group now appears to be inactive [^1] https://lists.w3.org/Archives/Public/public-rsp/2023Mar/$0$.html.
Time-Annotated RDF (TA-RDF) [48] is an alternative RDF representation for the timestamped stream problem (more precisely, TA-RDF has a direct translation to RDF). It identifies each stream element as an URI node in an RDF graph. The data relevant to that stream element is then attached to the node. This boils down to each stream element being an RDF graph in which a single node serves as the “point of entry”. This model is notably less flexible than the RSP Data Model. A similar approach is used by Stream Containers [49], a framework for publishing RDF streams on the Web. There, a stream is a sequence of unnamed RDF graphs that are referenced by their URIs. Each URI points to a node (the subject) within one of the graphs, and the temporal information is also contained within this graph.
One auxiliary RDF stream definition also appears in this category. The VoCaLS vocabulary for describing linked streams [59] defines an RDF stream as a potentially infinite sequence of RDF graphs and/or triples. This definition is used in practice in the RDF4J API [57], whose YASPER implementation can consume streams of RDF graphs or RDF triples.
2.3 Streaming Semantic Annotation and Translation
Several systems were developed for streaming semantic annotation (lifting) and translation. RMLStreamer [24, 45] annotates data using the RML language [15], outputting a stream of RDF datasets. CARML [9] is a Java library that provides generic facilities for annotating data using RML. It supports outputting a stream of RDF quad statements (using RDF4J Rio [16]), which corresponds to a single RDF dataset. CARML does not natively support streams of datasets, and thus it operates on a different abstraction level than RMLStreamer. This limitation is addressed by the Semantic Annotation enabler [55] from ASSIST-IoT [54], which outputs a stream of RDF datasets. SPARQL-Generate [38] is an extension of the SPARQL language, which allows for generating RDF from various data sources. Its reference implementation [^2] https://github.com/sparql-generate/sparql-generate supports outputting a stream of RDF triples (using Apache Jena RIOT [2]).
Last in this category is the Inter Platform Semantic Mediator (IPSM) [20], which translates a stream of RDF graphs into another such stream, in real time. Although on a logical level IPSM processes RDF graphs, it outputs a stream of datasets, where each dataset has two named graphs – one containing the translated RDF, and the other with metadata about the translation itself.
2.4 Semantic Streaming Applications
Over the years, RDF streams were employed in various real-life use cases. One such example is DBpedia-Live [44] – a now-defunct service that published changes to the DBpedia knowledge base in real time. Two streams were published – deleted and added triples. In each stream, the elements were RDF graphs, in the form of N-Triples files.
RDF streams are also commonly used in semantic sensor networks [39]. In this context, the Graph of Things (GoT) [36] was proposed, which uses streams of RDF graphs. Another common use case is monitoring social media activity. Here, a survey by Keskisärkkä & Blomqvist [29] noted that RDF streams should have RDF graphs as elements. However, they also acknowledge that RDF streams can have RDF triples as elements. The recent proposal for a live open scientific knowledge graph [37] uses streams of RDF graphs to monitor scientific and social activity in real time.
2.5 Streaming I/O
The last broad category of RDF streaming solutions are those that process simple sequences of RDF statements (triples or quads) for input/output (I/O) operations. Two examples are the RDF4J Rio toolkit [16] and Apache Jena’s RIOT [2], both of which can parse and serialize streams of statements, reading from or writing to a Java byte stream. The two most obvious serializations for this task are N-Triples and N-Quads, however, Jena also supports streamed block formats for Turtle and Trig, where statements are transmitted in small batches (blocks).
2.6 Summary
The review uncovered a wide variety of task formulations for RDF streams. It should be noted that most of the identified works use the same term (RDF stream) to describe very different concepts. The review also highlighted an important distinction between theoretical models of RDF streams (e.g., timestamped triples) and practical ones. This distinction was also observed earlier in several works [5, 8, 12, 49]. Another contrasting feature is the use of graphs or datasets as stream elements, versus triple or quad statements. This was also noticed in earlier works focusing on the practical aspects of implementing RDF streams [5, 29].
Worth noting is that some works have several views of the same RDF stream. For example, ERI [17] views its RDF streams as a flow of triples, blocks (RDF graphs), or subject-molecules (RDF subgraphs), depending on the level of abstraction. This suggests that describing the type of an RDF stream depends on the context in which it is considered, or, metaphorically, one’s perspective. In fact, an N-Triples file would be considered by RDF4J’s Rio [16] as a stream of triples, while an application like DBpedia-Live [44] or IPSM [20] would see it as a single element in a stream of graphs.
3 Proposed Systematization
In what follows, we propose a set of definitions for the types of RDF streams present in the literature and current practice. Then, Section 3.2 organizes these definitions in the RDF Stream Taxonomy, and Section 3.3 links the proposed stream types to the state of the art. The basis for all following definitions is the RDF 1.1 W3C Recommendation [10], with the aim of making the definitions readily applicable in the context of existing, RDF 1.1-compliant software.
The model-theoric semantics of RDF streams are not considered in this work. Instead, the proposed semi-formal definitions can be used as a universal basis on top of which the formal semantics can be defined, as needed. This is motivated by, firstly, the semantics of RDF streams being irrelevant for some of the discussed application areas (e.g., streaming protocols, streaming I/O). Secondly, only the semantics of RDF graphs are well-defined – there are no agreed-upon formal semantics for RDF datasets [66], and there are similarly many ways to approach this topic for RDF streams. Therefore, we leave this for future discussion.
3.1 RDF Stream Definitions
We start by defining a universal, abstract RDF stream that will serve as a common denominator for further discussion:
Definition 1
An RDF stream is an ordered, potentially infinite sequence of RDF stream elements.
This definition by itself cannot be used in practice, as it does not specify what are the RDF stream elements. Thus, we specialize this definition further:
Definition 2
A grouped RDF stream is an RDF stream whose elements are either RDF graphs or RDF datasets.
Definition 3
A flat RDF stream is an RDF stream whose elements are statements (either RDF triples or RDF quads).
Grouped and flat RDF streams are still abstract, however, they correspond to an important distinction between streams of groups of statements and streams of statements. With this, we can define the concrete types of RDF streams:
Definition 4
An RDF graph stream is a grouped RDF stream whose elements are unnamed (default) RDF graphs.
In other words, an element of an RDF graph stream is a set of triples. This is a commonly used formulation that can be further specialized:
Definition 5
An RDF subject graph stream is an RDF graph stream in which every element contains an IRI node (called the subject node) that uniquely identifies the graph in the stream. Every other node in the graph can be reached by traversing triples, starting from the subject node.
RDF subject graph streams or similar formulations were proposed to handle timestamped RDF streams (TA-RDF [48] and Stream Containers [49]). ERI [17] uses a similar formulation for subject-molecules, but its subject IRIs are unique only within one block. The proposed definition is slightly broader than the timestamped models, while still being useful – the advantage is that the element can be identified by an IRI, without using RDF datasets. The rest of the element can be discovered with what can be intuitively thought of as RDF graph connectivity.
We can then move on to streams of RDF datasets:
Definition 6
An RDF dataset stream is a grouped RDF stream whose elements are RDF datasets.
It should be noted here that although an RDF graph stream may seem like a special case of an RDF dataset stream, it is not. According to RDF 1.1, an RDF dataset is not a generalization of an RDF graph, but rather a collection of RDF graphs. Thus, the definitions of RDF dataset streams and RDF graph streams are not directly related in the taxonomy. It should be noted, however, that an RDF graph stream can be trivially transformed into an RDF dataset stream.
This definition can be further restricted to one named graph per element:
Definition 7
An RDF named graph stream is an RDF dataset stream in which every element has exactly one named RDF graph pair $\langle n,G\rangle$, where $G$ is an RDF graph, and $n$ is the graph name. Apart from graph $G$, the dataset may contain any number of triples in the default graph.
This is a simplified version of the RSP Data Model [47], without the temporal aspect. Although we could find only one reference to this stream type in the literature (RSP engines can consume such streams [47]), it is an intuitively important specialization of the RDF dataset stream. Here, every element corresponds to one named RDF graph with some optional information (metadata) about it. As with RDF subject graph streams, this has the advantage of identifying an element by IRI, albeit with more flexibility as to the element’s contents.
With this, we can move on to the timestamped variant:
Definition 8
A timestamped named graph is an RDF dataset in which:
(1) there is exactly one named RDF graph pair $\langle n,G\rangle$, where $G$ is an RDF graph, and $n$ is the graph name;
(2) the default graph includes a timestamp triple $\langle n,p,t\rangle$, where $p$ is a timestamp predicate that relates $t$, called the timestamp, and the graph $G$.
Definition 9
A timestamped RDF named graph stream is an RDF named graph stream in which every element is a timestamped named graph. The elements that share the same timestamp predicate $p$ are ordered by the partial order associated with $p$.
These definitions were derived directly from the draft RSP Data Model [47].
Finally, we define concrete flat RDF stream types:
Definition 10
A flat RDF triple stream is a flat RDF stream whose elements are triples.
Definition 11
A flat RDF quad stream is a flat RDF stream whose elements are quads.
An “RDF quad” is a term that is often used by practitioners, but not defined in the RDF 1.1 Recommendation. It is however defined in the RDF 1.2 Working Draft [25], and thus we use the RDF 1.2 definition of quads here.
Flat RDF streams can be represented simply as N-Triples or N-Quads files, however, contrary to RDF graphs and datasets, they have a defined order of statements. Similarly to RDF graph and dataset streams, a triple stream is not a subclass of a quad stream, because a triple is not a special case of a quad. However, a trivial transformation from one to the other can be performed, by stating explicitly that the triples belong to the default graph.
Grouped and flat RDF streams are notably interrelated, as any RDF graph stream can be flattened into an RDF triple stream, by simply listing and concatenating the statements in each of its elements – the same is true for RDF dataset streams and RDF quad streams. The reverse is also possible, the elements of a flat stream can be grouped into a grouped stream. Such operations are in fact common in streaming applications [5, 17, 29, 50].
Finally, for all of the above definitions, their generalized variants can be trivially derived (e.g., generalized RDF graph stream), by using generalized triples and generalized datasets. We skip these definitions for the sake of brevity.
3.2 RDF Stream Taxonomy
The proposed definitions form a taxonomical structure of task formulations (RDF Stream Taxonomy – RDF-STaX), summarized in Figure 1. This overview allows us to observe that each concrete RDF stream type uses one of the four element types: graph, dataset, triple, and quad, possibly with additional restrictions on them. In fact, these four types of elements seem to be the only conceivable ones that can be derived from the RDF 1.1 specification. For each of them, representing the element is a known and solved problem (RDF serializations), and thus they can be easily used in practical applications.
Figure 1: Overview of the RDF Stream Taxonomy (RDF-STaX).
3.3 Correspondence to the State of the Art
Finally, we link the proposed RDF stream types to the definitions used in state-of-the-art research and software – Table 1 presents such a summary. As noted in Section 2.6, a single RDF stream can be viewed from multiple perspectives, which is reflected in the table, with some items appearing more than once. In some cases the correspondence is not ideal and there may be insignificant differences between the definitions proposed here and those in the literature. The short descriptions presented in the table are expanded upon in the nanopublications dataset, as described in Section 5.1.
Stream type Uses
RDF graph stream
DBpedia-Live [44], ERI (blocks) [17], GoT [36], IPSM (logical level) [20], Jelly (stream frames) [50], Keskisärkkä & Blomqvist [29], Live open scientific knowledge graphs [37], RDF EXI [28], RDSZ [18], RSP Data Model (can be consumed) [47], RSP4J YASPER (can be consumed) [57], S-HDT [26], VoCaLS [59]
RDF subject
graph stream
ERI (subject-molecule stream, only within a block) [17], TA-RDF [48], Stream Containers [49]
RDF dataset stream
IPSM (output) [20], Jelly (stream frames) [50], RMLStreamer [24, 45], RSP Data Model (can be consumed) [47], Semantic Annotation enabler [55]
RDF named
graph stream
RSP Data Model (can be consumed) [47]
Timestamped RDF
named graph stream
Linked Data Notifications for RDF streams [7], RSP Data Model [47], Tappolet & Bernstein [56], TripleWave [41], Wu et al. [63]
Flat RDF
triple stream
Apache Jena RIOT [2], ERI (high level) [17], Groppe et al. [23], Jelly (high level) [50], Keskisärkkä & Blomqvist (mentioned) [29], RDF4J Rio [16], RSP4J YASPER (can be consumed) [57], SPARQL-Generate [38], VoCaLS [59]
Flat RDF
quad stream
Apache Jena RIOT [2], CARML [9], Jelly (high level) [50], RDF4J Rio [16]
Table 1: RDF stream type usage in research and software, according to RDF-STaX.
Different types of RDF streams appear to be applicable in different settings. In the context of network streaming applications, most commonly used are grouped RDF streams. Flat RDF streams, on the other hand, are used for streaming I/O and internally by various applications.
4 RDF-STaX Ontology
To make the taxonomy readily applicable, it was modeled as an OWL 2 DL ontology [^3] https://w3id.org/stax/1.0.1/ontology (Figure 2). Stream types are instances of either stax:AbstractStreamType or stax:ConcreteStreamType class, with taxonomical relations being realized using SKOS in the OWL DL version [27, 42]. Each stream type is accompanied by a label, formal definition, and usage example.
Figure 2: Classes and object properties in the RDF-STaX ontology.
The ontology includes rich semantic relations between stream types. Properties stax:canBeFlattenedInto and stax:canBeGroupedInto relate grouped and flat stream types that can be easily converted into one another (e.g., an RDF graph stream can be flattened into a flat RDF triple stream). The property stax:canBeTriviallyExtendedInto is used for graph- and triple-based streams, to indicate that they can be trivially turned into dataset- and quad-based streams, respectively. This transformation consists of stating explicitly that the triples are in the default graph. Object property chains are defined in the ontology to allow reasoning on these relations, exploiting the taxonomical relations from SKOS.
It should be stressed that RDF-STaX does not define a class for RDF streams, but rather classes for RDF stream types. It also defines a class for RDF stream type usages, which can be thought of as subjective views of how an RDF stream is used. As described in detail in the documentation [^4] https://w3id.org/stax/use-it, the expected usage pattern is for users to create instances of the stax:RdfStreamTypeUsage class to annotate research works, software, or datasets, with each instance corresponding to a different view or use case of the stream. This pattern was selected due to the inherent subjectivity of determining the types of RDF streams, as discussed in Section 2.6. Below is an example of how this pattern may be used to annotate a DCAT [6] Dataset:
@prefix dcat: <http://www.w3.org/ns/dcat#> .
@prefix rdfs: <http://www.w3.org/2000/01/rdf-schema#> .
@prefix stax: <https://w3id.org/stax/ontology#> .
[…]
#  other properties of the dataset
stax:hasStreamTypeUsage [
a stax:RdfStreamTypeUsage ;
stax:hasStreamType stax:datasetStream ;
rdfs:comment "The data is a sequence of RDF datasets."@en
%**** main.tex Line 350 ****a stax:RdfStreamTypeUsage ;
stax:hasStreamType stax:flatQuadStream ;
rdfs:comment "The data can be viewed as a flat sequence of RDF quads."@en
The ontology’s documentationfootnote 3 contains a detailed description of every class, property, and individual.
4.1 Availability and Sustainability
The sources of the ontology and its documentation are hosted on GitHub [^5] https://github.com/RDF-STaX/rdf-stax.github.io. The repository includes continuous integration and deployment (CI/CD) scripts that automatically validate the ontology’s profile, generate documentation, run an OWL 2 reasoner to add more assertions, package the ontology into a number of formats, and publish it on the website [^6] https://w3id.org/stax. Two versions of the ontology are published: OWL 2 DL, which can be used with reasoners; and OWL 2 Full, which contains more metadata and alignments to other vocabularies.
The ontology uses best findability, accessibility, interoperability, and reusability (FAIR) practices developed by the community [22], achieving 86% score in FOOPS! [21] and 4/4 stars in DBpedia Archivo [19]. All releases of the ontology, including the OWL 2 DL and OWL 2 Full variants, are distributed through the RDF-STaX website, under a permanent URL. The ontology is also archived in Zenodo [51] and registered in DBpedia Archivo and Linked Open Vocabularies [61] (pending acceptance).
RDF-STaX is an open and freely-licensed project, welcoming contributions to the ontology, its documentation, and tooling. A usage guidefootnote 4 with several use case scenarios and examples was prepared to facilitate adoption. To ensure long-term sustainability, a comprehensive contribution guide was provided [^7] https://w3id.org/stax/1.0.1/contributing, along with a public issue tracker [^8] https://github.com/RDF-STaX/rdf-stax.github.io/issues. The project is entirely hosted on permanently free infrastructure, which should ensure its long-term stability.
4.2 Comparison with Other Ontologies
To the best of our knowledge, there is no other ontology or vocabulary that would model RDF stream types. As previously noted, the RDF-STaX ontology purposefully does not define classes for RDF streams, but rather focuses on subjective RDF stream usages. This difference makes other ontologies non-overlapping – instead, RDF-STaX can naturally complement them. To this end, the documentation includes examples of using RDF-STaX with these complementary ontologies (VoCaLS, VoID, and DCAT)footnote 4. The usage pattern is to define an RDF stream / dataset with these ontologies as usual, and add the stax:hasStreamTypeUsage property to indicate the stream type, according to RDF-STaX. The same pattern can be used with any future ontology for describing streams or datasets.
The VoCaLS vocabulary for describing published RDF streams [59] defines the RDFStream class for potentially infinite sequences of RDF graphs and/or triples. These streams can be thus annotated as RDF graph streams or flat RDF triple streams in RDF-STaX. The VoID vocabulary [65] describes datasets with RDF data (not RDF datasets as per RDF 1.1), which can be viewed as flat RDF triple streams. The more general DCAT [6] describes datasets and their distributions, which may be RDF streams.
The RDF-STaX ontology also defines four element types (RDF graphs, triples, quads, and datasets). In this regard, we could identify only the SPARQL 1.1 Service Description vocabulary [62], which has the sd:Graph and sd:Dataset classes for RDF graphs and datasets. These classes were aligned with the corresponding terms in RDF-STaX, using the skos:relatedMatch property.
5 Applications
RDF-STaX was created with several applications in mind. Firstly, it can be used to describe stream types used in research and software, to easily compare different solutions and foster compatibility. Secondly, it may be employed to semantically annotate published datasets and streams, improving their accessibility and interoperability. Both of these use cases have already been realized in practice.
5.1 Nanopublications
A nanopublication is a small unit of scientific knowledge published in RDF [31] – it can be an opinion, a measurement, or any other assertion. Nanopublications are a natural way to express statements about scientific works or software due to their strong provenance mechanisms and citeability. This fits well the pattern of subjective statements about stream types, which is used in RDF-STaX. To make it easier to create nanopublications with the RDF-STaX ontology, a template was prepared for Nanodash, an open nanopublication editor [32]. The template along with the accompanying manual [^9] https://w3id.org/stax/1.0.1/nanopubs makes it easy to create and publish semantic assertions such as “X uses stream type Y because Z”, requiring the user only to fill out a simple online form.
The results from the survey conducted in this paper (as summarized in Section 3.3) were published as 32 nanopublications, including a more detailed discussion of the usages than what could fit in the paper. The RDF-STaX nanopublications are packaged as a single knowledge graph in a CI/CD pipeline and published automatically on the websitefootnote 9. Anyone can contribute to the dataset by simply publishing nanopublications with the provided template. The dataset can serve as a living state-of-the-art review, allowing one to easily compare and discuss works on RDF streaming, in a structured manner.
5.2 RiverBench
RiverBench is an open RDF streaming benchmark suite, containing heterogeneous streaming datasets that are published using automated pipelines, following FAIR principles [52]. RiverBench integrates with the RDF-STaX ontology, requiring all datasets to be annotated with their stream types. Each dataset has two stream types assigned, one for the grouped stream formulation and another for the flat. The consistency of the annotations is checked in CI/CD using SHACL rules and the semantic relations defined in RDF-STaX (SKOS taxonomy properties and stax:canBeFlattenedInto). The datasets are then organized into benchmark profiles by their stream type. The published dataset documentation and RDF metadata based on DCAT [6] also includes the RDF stream types.
At the time of writing, RiverBench’s 12 datasets use each stream type in RDF-STaX, except the RDF named graph stream (but there are examples of the timestamped variant). Detailed information about this use case along with examples are available on the website [^10] https://w3id.org/stax/1.0.1/uses.
6 Discussion and Future Work
Although the “RDF stream” is a term that is now ubiquitous in research, it is used to describe many different concepts. There is no universally accepted language for describing RDF streams, which leads to a situation in which RDF streaming solutions become hard to assess, compare, and connect together. For example, forming a pipeline composed of a stream annotator, translator, streaming protocol, and a reasoner currently appears to be a tough challenge, as it is hard to determine what exactly is an RDF stream for each of these tools.
It appears that the variety of RDF stream formulations is here to stay. Each of the different definitions has its merits and use cases, and none of them should be disregarded. Instead, we call to embrace the variety of RDF streams by systematizing them in one shared taxonomy. The taxonomy with its semantic relations (taxonomical, flattening, grouping, extending – see Section 4) allows one to reason about stream types not only in terms of subclassing, but also in terms of possible conversions between them. This, we hope, will be a much needed step in the direction of making RDF stream applications more compatible with each other.
RDF-STaX is embodied in an ontology that makes the taxonomy readily usable in practice. The ontology does not attempt to replace any existing vocabulary for describing streams or datasets, but rather aims to complement them. We hope that by being non-overlapping by design and focused on a single purpose, RDF-STaX will remain useful and relevant in the long term. The two presented use cases, both of which were already realized in practice, are very common and well-known to the community, making the ontology all the more relevant.
Another result of this work is the nanopublications dataset, which is a collaborative, living state-of-the-art review of RDF streaming. Nanopublications, in our view, are a natural way to structure intersubjective scientific discourse. It is possible for example, to comment on other nanopublications, confirm or disagree with them, all in a structured manner. Here we propose to use this powerful tool for a very narrow field of science, but its possible future applications are certainly much broader. We encourage other researchers to contribute their perspectives to the dataset or to comment on the existing assertions.
As for future work, the theory presented in this contribution notably does not tackle the formal semantics of RDF streams. As with RDF datasets, it is unlikely that there can be single, unifying semantics for streams, but nonetheless, we see this area as worth exploring in the future. The presented survey of RDF streaming is also not a systematic review – it only serves as a basis for the discussion in this work. Producing a systematic review of RDF streaming would be very valuable to the community, and we hope that the proposed taxonomy and nanopublication tooling will make creating it easier.
7 Conclusion
In this work we propose the RDF Stream Taxonomy, which systematizes RDF stream definitions found in the literature. The taxonomy uncovered the variety of RDF streaming in research and practice, highlighting the open challenge of making the different applications compatible with each other. The proposed RDF-STaX ontology embodies the semantic relations between stream types in the taxonomy, and makes the definitions readily applicable. The ontology positions itself as complementary to existing vocabularies, and employs a flexible approach of making subjective statements about the types of RDF streams. The ontology and its documentation are fully open and follow best practices for FAIR developed by the community.
The ontology was applied in two use cases – annotating research works in nanopublications, and describing RDF stream types of streaming datasets. The nanopublications dataset is a living state-of-the-art review that anyone can contribute to, using the provided documentation and tooling. The second use case highlights how RDF-STaX can be used to annotate published streaming datasets, integrating with other vocabularies.